X^2+16X+65=(X+7)*(_____)+(2)
X^2+10X+27=(X+6)*(_____)+(3)

Answer by xcentaur(357)

Using synthetic division in both cases,
X^2+16X+65=(X+7)*(_____)+(2)

x+7 | X^2 + 16X + 65
 -7 | 1     16    65
    | -     -7    -63
    ----------------
      1     9     2

Ans: (x+7)(x+9)+2


X^2+10X+27=(X+6)*(_____)+(3)


x+6 | X^2 + 10X + 27
 -6 |  1    10    27
    |  -    -6    -24
    -----------------
       1     4    3

Ans: (x+6)(x+4)+3


Hope this helps,
Prabhat